ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

1991 MS Classification: 0 Mori, Masao. Chuo University, Japan Beginners Confusion and Terminology in Mathematics Things have not much improved since the author reported at ICM 90 that most Japanese high school students failed to understand the necessary condition. In mathematics, it is defined as an inevitable result, however in daily usage, it means an indispensable precondition. ’Result’ and ’precondition’ are logically opposite to each other. We can hardly say that we have given the beginners sufficient and proper explanation about our peculiar terminology. any beginners should feel confusion and possibly hate toward a part of mathematics as before. We have been disturbing their logical sense, overlooking their difficulty and failing to manage it. This kind of educational problems do not seem native only to Japan. -MORI Masao Department of Mathematics, Chuo University 1-13-27 Kasuga Bunkyoku, Tokyo 112 Japan